The IBM nine-track tapes that became the industry standard for storage for three decades had several sizes , the most common being 2400 feet. The high density version stored data at 1600 bits/inch/track. If the whole tape (not including the parity track) could be used for data storage , how much raw data could one of these tapes hold?

Round to the nearest MB using decimal MB not binary MB
Answer #1

1 foot = 12 inches

2400/9=266.66 feet per track

1 track excluding, remaining 8 track =2400 - 266.66 = 2133.33 feet

so, space available to store data = 2133.33 feet

now, 2133.33 feet = 25600 inches

total data that can be stored = 25600 * 1600 = 40960000 bits

= 46080000/8 = 5120000 Bytes

= 5120000/1000000 = 5.12 MB = 5 MB (approx.)